Calculate the standard molar entropy change for each of the following reactions at 25 oC.

  1. Ca (s) + 2H2O (l) → Ca(OH)(aq) + 2H(g)
  2. Na2CO(aq) + 2HCl (aq) → 2NaCl (aq) + H2O (l) + CO(g)
1
Expert's answer
2021-11-19T14:15:03-0500

Solution:

The change in the standard molar entropy of a reaction can be found by the difference between the sum of the molar entropies of the products and the sum of the molar entropies of the reactants.

ΔS°reaction = Σnpproducts - Σnrreactants


(1):

Ca (s) + 2H2O (l) → Ca(OH)(aq) + H(g)


S°(Ca, s) = 41.42 J mol-1 K-1

S°(H2O, l) = 69.91 J mol-1 K-1

S°(Ca(OH)2, aq) = -74.5 J mol-1 K-1

S°(H2, g) = 130.684 J mol-1 K-1

Hence,

ΔS°reaction = (S°Ca(OH)2 + S°H2) - (S°Ca + 2×S°H2O)

ΔS°reaction = (-74.5 + 130.684) - (41.42 + 2×69.91) = -125.056

ΔS°reaction = -125.06 J mol-1 K-1


Answer (1): ΔS°reaction = -125.06 J mol-1 K-1


(2):

Na2CO(aq) + 2HCl(aq) → 2NaCl (aq) + H2O (l) + CO(g)


S°(Na2CO3, aq) = 61.1 J mol-1 K-1

S°(HCl, aq) = 56.5 J mol-1 K-1

S°(NaCl, aq) = 115.5 J mol-1 K-1

S°(H2O, l) = 69.91 J mol-1 K-1

S°(CO2, g) = 213.74 J mol-1 K-1

Hence,

ΔS°reaction = (2×S°NaCl + H2O + S°CO2) - (S°Na2CO3 + 2×S°HCl)

ΔS°reaction = (2×115.5 + 69.91 + 213.74) - (61.1 + 2×56.5) = 340.55

ΔS°reaction = 340.55 J mol-1 K-1


Answer (2): ΔS°reaction = 340.55 J mol-1 K-1